OpenStudy (anonymous):

3(x - 2) + 4(2x - 6) = 6(x-4) + 8(2x + 1) X=?

OpenStudy (anonymous):

3x-6+8x-24=6x-24+16x+8 11x-30=22x-16 -11x=14 x=-14/11

OpenStudy (anonymous):

Aditi's right..